OpenStudy (nsh4267):

If you earn $2900 per month and you expect your earnings to increase by 3.8% per year, how much do you think you will be making in 8 years? (Express your answer rounded correctly to the nearest cent!)

OpenStudy (wolf1728):

Well 2,900 per month equals $34,800 per year. Salary = 34,800 * (1.038)^n where n = # of years After 8 years Salary = 34,800 * (1.038)^8 Salary = 34,800 * 1.3476553139 Salary = 46,898.40

OpenStudy (hermom1967):

And then divide $46,898.40 by 12 months to get your monthly salary

OpenStudy (wolf1728):

Actually, it says "how much do you think you'll be making in 8 years" - It doesn't state whether it's monthly or annually.